1. An aqueous waste stream in a metal finishing plant has a flow rate of 1000 gal/d and contains 100 ppm of ionic copper (i.e., Cu2+). The waste stream has a very low pH of 0.3 because concentrated nitric acid is used in one of the etching processes. Engineers plan to treat the waste stream with slaked lime and raise the pH up to circumneutral.

(a) Assuming nitric acid completely dissociates to H+ under these conditions, what concentration of HNO3 (in moles per liter) would be required to achieve pH 0.3 in the waste stream above. Assume ideal solution conditions prevail.

(b) Given that nitric acid has pKa ≈ –1, what concentration of HNO3 (in moles per liter) is actually required to achieve pH 0.3 in the waste stream above. Once again, assume ideal solution conditions prevail. (c) Based on solubility considerations, how much of the copper might be removed as a precipitate during neutralization of the waste stream above? Assume the pH is raised from 0.3 to 7 during neutralization and that the solubility product for this reaction is: Ksp = [Cu2+] [OH-]^2 = 1.6 ×10^-19

(d) Taking into account the reactions in parts (b) and (c) above, determine the total amount of Ca(OH)2 (in lbm per day) that will be consumed upon neutralization of the waste stream and subsequent copper precipitation.
Answer #1

[H+] = 1-0, 3(M) = 0.50 (M) oa endu | dissociates to ttdirns Ini Hal Mbl Cod (Molarity) Cot.CO 2. Ka1 or log ka or Ka, 10-10New, putting d.е%) into equation@ we get o /20 Cx29-0.50 2l 2.0 O. 52501) The aetual cencoutvelion of Nog rquid 525 male bH sTotelamoud of waste shrem per day 1000 al 1000x 322 5.6 32256 = 32256 x 106mL 3 3225 . 6 x 1-5598 1рґ mol = 3225-6 x 1.55980